10. Sketch the graph of -5x2- 16xy +7y2-198 0. Show the steps used in rotating the axes
Answer #1

Date Page -5と-16Xyt7y=198 -.G) T-s -8 1 dktmtij deoveov xtAx19-87( engen vame A dvn det(A AT) D 5- - -8 -64=0 -35 -2λt N, A,d. -8C +7d d -2C iran-ik es vedtar far , the M. new G 24 2M 24 S hiieitaonta htan the es anned with CamomomaDate Paga 3 nsm Gi) amd (i we PY)A(eY) t8 9 0 C ylaabeta th mojen X Systena transformed Scanned With Cam Scanner